OpenStudy (anonymous):

Carla earns $8.43 per hour working 40 hours a week. If she worked 7 overtime hours one week, what would be her earnings?

OpenStudy (mertsj):

Does she get time and a half for overtime?

jimthompson5910 (jim_thompson5910):

She worked 40 hours at $8.43, so she earned 40*8.43 = 337.20 dollars Then she works an additional 7 hours at a wage that's 1.5 times the normal wage. So she worked 7 hours at 1.5*8.43 = 12.65 dollars an hour to give her total overtime earnings to be 7*12.65 = 88.55 Now just add up the individual totals: 337.20+88.55 = 425.75 So her total earnings is $425.75